HUBBLE SPACE TELESCOPE DAILY REPORT #2229 PERIOD COVERED: 0000Z (UTC) 10/14/98 - 0000Z (UTC) 10/15/98 Daily Status Report as of 288/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. As described in yesterday's report and HSTAR 6714, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were executed as scheduled with no further anomalies. 1.2 Completed Two Sets of WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. As described in yesterday's report and HSTAR 6714, the acquisition for the first iteration of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the procedures were conducted as scheduled, and no further reported problems. 1.3 Completed Two Sets of NIC/2 7441 (A Search for Zodiacal Dust around Bright Nearby Stars) The Near Infrared Camera (Camera Two) was used to make observations of the zodiacal lights around the bright nearby star Altair. Dust particles around a main sequence star are quickly removed by Poynting-Robertson drag, therefore, the existence of such dust is an almost certain signature of a current presence of a population of larger bodies in orbit around the star. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Four Sets of NIC/1/2/3 7963 (DARKs: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to >100K. The observations completed with and no anomalies. 1.5 Completed NIC/2 7897 (Probing Planetary Formation Around Main-Sequence Stars: A Snapshot Survey) The Near Infrared Camera (Camera Two) was used to make snapshort survey images of the main sequence star HR123 probing for planetary formation. The IR excess main-sequence stars host cool circumstellar dust disks, which are most likely replenished by the debris of unseen planetesimals. The observations were completed with no problems. 1.6 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star LHS416. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.7 Completed NIC/1/2/3 7886 (NICMOS Snap Shot Survey of Early-Type Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the early-type galaxy NGC-1426 as part of a snap-shot broad-band imaging survey. Since almost all early-type galaxies harbor dust, these images which are less sensitive to extinction and will provide superior light profiles for dynamical studies of the cores of these galaxies. The observations were completed as scheduled, and no problems were reported. 1.8 Completed FGS 7491 (The Masses of the Brown Dwarf Candidate GL 623B) The FGSs were used to conduct astrometry at 1-mas accuracy of faint companions to nearby M-dwarfs. This is the most productive direct method to determine accurate masses of stars near the end of the main sequence. From preliminary FGS results we estimate the mass of W922 as 0.10 solar masses, placing both near the stellar/substellar border. Such a mass determination for W922 should establish conclusively its nature as star or substellar object. The observations were executed as scheduled with no anomalies. 1.9 Completed WF/PC-2 7386 (A Snapshot Survey of Nova Shells) The WF/PC-2 was used to make observational images of the super nova shell V2290-OPH in order to detect and measure its expanding gaseous shell. These observations will trace the distribution of hot and cool gas in the nova shells, and show their degree of inhomogeneity on both large and small scales. The observations were completed as scheduled, and no problems were noted. 1.10 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.11 Completed NIC/1/2/3 7893 (The Two Types of BL Lac Objects: Extrema of Jet Physics) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the BL Lac object H2356-309. These observations will build on the WF/PC-2 survey of a well-defined sample of 97 BL Lac objects, for which host galaxies were detected in 90 simulations predict ~100 provide essential constraints on galaxy properties -- luminosity, size, morphology, and color -- including an unprecedented ensemble of host galaxy color profiles at subarcsecond spatial resolution. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2STIS/CCD/NIC/1/2/3 8071 (HDF-South Flanking Fields) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD) and the Near Infrared Camera (Cameras One, Two and Three) were all used to make observations of the distant object SAO255267 as part of the HDF-South flanking field study. This program is designed to provide coverage around the HDF-South field and to provide a deep optical images of the NICMOS HDF-South field. The images were completed as planned, and no problems were reported. 1.13 Completed WF/PC-2 7623 (WF/PC-2 Cycle 7 Internal Flats) The WF/PC-2 was used to make a set of internal Cycle 7 calibration flats which are designed to monitor the pixel to pixel flat field response and the visflat lamp degradation, as well as detect any possible changes due to contamination. The calibration observations were completed as planned, and no problems were reported. 1.14 Completed STIS/CCD/MA1/MA2 7673 (STIS MAMA Sensitivity and Focus Monitor) The Space Telescope Imaging Spectrograph (CCD/MA1/MA2) was used to make observations of standard star GRW+70D582 in order to perform a MAMA sensitivity and focus monitoring program. This is to monitor sensitivity of each MAMA grating mode to detect any change due to contamination or other causes. A wavelength calibration image was also taken. The observations were executed as scheduled with no anomalies. 1.15 Completed NIC/1/2/3 7869 (The Morphology of Dense Gas in Seyferts , Obstruction and Fueling of AGNs) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the Seyfert galaxy NGC-2992 to study the morphology of its dense gas. Near-Infrared color maps and images in H-2 emission will be used to trace hot and cold dense gas in this Seyfert galaxy. The observations were executed as scheduled with no anomalies. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Per HSTAR 6716, the acquisition at 288/045637Z defaulted to fine lock backup on FGS-1. Proposals appearing in tomorrow's report (WF/PC-2 6703 and NICMOS 8082) may have been affected. Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: The SSR EDAC error counter was cleared twice, using ROP SR-1. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH